Megatronics Corporation, a massive retailer of electronic products, is organized in four separate divisions. The four divisional managers are evaluated at year-end, and bonuses are awarded based on ROI. Last year, the company as a whole produced a 13 percent return on its investment.

During the past week, management of the company’s Northeast Division was approached about the possibility of buying a competitor that had decided to redirect its retail activities. (If the competitor is acquired, it will be acquired at its book value.) The data that follow relate to recent performance of the Northeast Division and the competitor:

Northeast Division Competitor
Sales $ 4,260,000 $ 2,660,000
Variable costs 70 % of sales 65 % of sales
Fixed costs $ 1,087,000 $ 859,000
Invested capital $ 955,000 $ 300,000

Management has determined that in order to upgrade the competitor to Megatronics’ standards, an additional $180,000 of invested capital would be needed.

Required:

  1. 1. Compute the current ROI of the Northeast Division and the division’s ROI if the competitor is acquired.

  2. 2. If divisional management is being evaluated on the basis of ROI, will the Northeast Division likely pursue acquisition of the competitor?

  3. 3-a. Compute the ROI of the competitor as it is now and after the intended upgrade.

  4. 3-b. If ROI is used as the basis for evaluation, would Megatronics Corporation likely be in favor of the acquisition of the competitor?

  5. 4. Calculate the Northeast Division's ROI after acquisition of competitor but before upgrading.

  6. 5-a. Assume that Megatronics uses residual income to evaluate performance and desires a 12 percent minimum return on invested capital. Compute the current residual income of the Northeast Division and the division’s residual income if the competitor is acquired.

  7. 5-b. If divisional management is being evaluated on the basis of residual income, will the Northeast Division likely pursue acquisition of the competitor?

Questions I need answered by chegg :)

Compute the current ROI of the Northeast Division and the division’s ROI if the competitor is acquired. (Round your answers to 2 decimal places (i.e., .1234 should be entered as 12.34).)

Current ROI %
ROI if competitor is acquired %

Compute the ROI of the competitor as it is now and after the intended upgrade.

ROI before upgrading %
ROI after upgrading

%

Calculate the Northeast Division's ROI after acquisition of competitor but before upgrading. (Round your answer to 2 decimal place. (i.e., .1234 should be entered as 12.34).)

ROI

%

Assume that Megatronics uses residual income to evaluate performance and desires a 12 percent minimum return on invested capital. Compute the current residual income of the Northeast Division and the division’s residual income if the competitor is acquired.

Current residual income
Residual income if competitor is acquired
Answer #1

1.ROI = (Sales – variable costs – Fixed costs)/Invested Capital

Current ROI = (4,260,000*30% - 1,087,000)/955,000

= 20%

If competitor is acquired = (191,000+ 72000)/1,435,000

= 18.33%

2.No, since it will reduce ROI

3-a Now = 72000/300,000 = 24%

After upgrade = 72000/480,000 = 15%

3-b. Yes, since ROI is higher than overall ROI

4.ROI = 263000/1,255,000

= 20.96%

5-a Current residual Income of Northeast division = Operating income – Invested capital*Required return

= 191,000 – 955,000*12%

= $76,400

If Acquired = 76,400 + 14400

= $90,800

Yes, since it will increase Residual Income